# Quotients of flag varieties and their birational geometry

### v1

@@ 1. Introduction
On the other hand, Kapranov, Białynicki-Birula, Sommese, et al. (see [21], [5]) considered the problem of constructing an intrinsic notion of quotients by group actions, and proposed the closely related concepts of Chow and Hilbert quotients of a variety.
@@
A remarkable point is that Chow quotients by the action of complex tori of complex rational homogeneous varieties – which are among the simplest and most symmetric algebraic varieties – provide extremely interesting examples of varieties. For instance, Kapranov showed that the Grothendieck–Knudsen moduli space of stable genus zero curves with marked points can be realized as the Chow quotient of a Grassmannian by the action of a maximal torus in its automorphism group (cf. [21, Theorem 4.1.8]). This stimulated other researchers, such as Thaddeus, who characterized certain classical moduli spaces – including the space of collineations – as Chow quotients (see [31, Main Theorem]).
A natural problem – that we address in this work – is to extend Kapranov’s study to Chow quotients of other rational homogeneous varieties by the action of a maximal torus.
Rational homogeneous varieties are defined as quotients of semisimple algebraic groups by parabolic subgroups, and occur in families, classified by the underlying semisimple group $G$. For a fixed $G$, all rational homogeneous $G$-varieties arise as smooth fiber type contractions of the associated complete flag manifold, which is the quotient of $G$ by a Borel subgroup $B$. This is the rational homogeneous variety that is “closest” to $G$, and in fact it has been shown that the homogeneous structure of $G/B$ can be reconstructed out of geometric properties of its elementary contractions (cf. [27], [28, Appendix A]). It makes then sense to start the research by looking at Chow quotients of complete flag varieties, from which the rest of quotients of rational homogeneous varieties can be obtained by means of contractions (see Section 8.3).
The geometry of these Chow quotients can be very involved, and substantially different from the varieties defining them. First of all, they can be singular, as one can check for the complete flag associated to the group $\mathop{\rm Sp}\nolimits(4)$ (cf. [6]). Then, while the geometry of a complete flag $F=G/B$ is determined by the action of the group $G$, the Chow quotient $X$ of $F$ is naturally endowed with the action of the Weyl group $W$ of $G$. Moreover, while all the contractions of $F$ are smooth fibrations, its quotient $X$ enjoys a frantic birational life.
One can easily show that the Chow quotient of the complete flag manifold associated to $\mathop{\rm PGL}\nolimits(3)$ is ${\mathbb{P}}^{1}$, hence the first interesting case is the complete flag variety of vector subspaces of ${\mathbb{C}}^{4}$. Our main result is the following:
@@
Let $F$ be the complete flag variety of vector subspaces of ${\mathbb{C}}^{4}$, endowed with the natural action of the complex torus $H\subset\mathop{\rm PGL}\nolimits(4)$ of homothety classes of diagonal matrices. Then the Chow quotient $X$ of $F$ by the action of $H$ is a smooth rational weak Fano threefold of Picard number $12$.
A large part of the paper is then devoted to the description of the birational properties of $X$. We will show that $X$ admits the action of an index two extension of the Weyl group, namely, the octahedral symmetry group $S_{4}\rtimes{\mathbb{Z}}/2{\mathbb{Z}}$. This action is then inherited by the space of numerical class of divisors ${\mathop{\rm N^{1}}\nolimits}(X)$, and by the nef and effective cones of $X$, which we will describe.
@@
We remark that in [13, Theorem 1.3] it was already shown, using other methods, that the Chow quotient $X$ is smooth (along with being a log crepant resolution of the log canonical compactification of an open subset of a subvariety of $F$ parametrizing flags with a fixed base point). Our approach is different, and we give an explicit description of $X$.
@@
The last two sections of the paper focus on the birational geometry of the variety $X$. After studying its intersection theory in terms of boundary divisors – which generate $\mathop{\rm Pic}\nolimits(X)$ – we prove that $X$ is a smooth weak Fano manifold, whose anticanonical model has degree $12$ in ${\mathbb{P}}^{13}$ (cf. Section 7.3); in particular this concludes the proof of Theorem 1.1.
As a consequence $X$ is a Mori Dream Space, therefore its Mori, Nef and Effective cones are polyhedral. We compute these cones in Section 8, and study some of the contractions of $X$. In particular, we identify the Chow quotients of partial flag varieties of ${\mathbb{C}}^{4}$ as targets of contractions of $X$, which we describe explicitly.
@@
Acknowledgements. We would like to thank J. Hong for presenting this problem to us, as well as J. Wiśniewski, M. Donten-Bury and E. Lee for interesting conversations and helpful references.
@@ 2.2. Torus actions and associated quotients
Consider again the above setting. Following [2, §2], we can describe a stronger relation among $\mathcal{C}\mkern-1.0muZ$ and the GIT quotients of $Z$. Let $u,v\in\Gamma$ be such that $Z^{ss}_{v}\subset Z^{ss}_{u}$. Taking quotients we obtain a dominant projective morphism $f_{v,u}:\mathcal{G}\!Z_{v}\to\mathcal{G}\!Z_{u}$. One may show that the collection of GIT quotients $\mathcal{G}\!Z_{v}$, for $v\in\Gamma$, together with the morphisms $f_{v,u}$, for $u\in\Gamma$ such that $Z^{ss}_{v}\subset Z^{ss}_{u}$, form an inverse system. The inverse limit of this system might be reducible (cf. [22, p. 654], [30, Counterexample 1.11]), but there exists a unique irreducible component $\overline{\mathcal{L}Z}$ containing the geometric quotient $\mathcal{G}U$. We call $\overline{\mathcal{L}Z}$ the limit quotient of $Z$, and its normalization $\mathcal{L}Z$ the normalized limit quotient of $Z$ by the action of $H$.
@@ 3.1. Fixed flags and their weights
See [26, Propositions 3.6, 3.7] and the references therein. ∎
@@ 4.5. Description of $X_{e}$ as an inverse limit of GIT quotients
  • The difference between these varieties is that the closed sets $\pi(Z_{(24)})\subset X(\Sigma_{+})$, $\pi(Z_{(03)})\subset X(\Sigma_{-})$ are contracted to the point $\pi(Z_{(0234)})$ in $X(\Sigma_{0})$. The common refinement of $\Sigma_{+}$, $\Sigma_{-}$, whose associated toric variety resolves the induced birational map $X(\Sigma_{+})\dashrightarrow X(\Sigma_{-})$, is the quotient fan $\Sigma$. One may check that the map $X(\Sigma_{+})\dashrightarrow X(\Sigma_{-})$ is locally the standard Atiyah flip (cf. [14, Example 11.1.12], [29, § 1.3]), and that we have a Cartesian square:
@@ 5.2. Generators of the group
We take a general element $x$ of $\mathfrak{n}$}, i.e., the class of a matrix
$Y=\begin{pmatrix}0&0&0&0\\ 1&0&0&0\\ y_{1}&1&0&0\\ y_{3}&y_{2}&1&0\\ \end{pmatrix}$,
which corresponds to the point $1:y_{1}:y_{2}:y_{3}]\in{\mathbb{P}}^{3}$.
We first compute the matrix $r_{i}e^{Y}$ (resp. $w_{0}((e^{Y})^{\top})^{-1}w_{0}$ in the case of the anti-transposition $\tau$, see the last item of Section 2.3), and its LU-decomposition.
We then define the matrix $M$ as the logarithm of the unipotent low triangular factor $L$, whose entries are rational functions in the variables $y_{i}$. We finally compute the class of $M$ modulo $H$, obtaining three rational functions $f_{1},f_{2},f_{3}$ in the variables $y_{i}$, interpret them as a rational map ${\mathbb{P}}^{3}\dashrightarrow{\mathbb{P}}^{3}$, sending $1:y_{1}:y_{2}:y_{3}]$ to $1:f_{1}:f_{2}:f_{3}]$, and homogenize these expressions, adding a new variable $y_{0}$.
We have performed the above procedure using SageMath, obtaining expressions for the birational maps associated to $r_{1},r_{2},r_{3},\tau$ in the homogeneous coordinates $y_{0}:y_{1}:y_{2}:y_{3}]$. These expressions take a particularly simple form after the following change of coordinates:
(5) $\begin{pmatrix}x_{0}\\ x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}6&0&0&0\\ 3&-6&0&0\\ 3&0&6&0\\ 2&-3&3&-6\end{pmatrix}\begin{pmatrix}y_{0}\\ y_{1}\\ y_{2}\\ y_{3}\end{pmatrix}$
The outcome we obtain is the following:
@@ 7.2. Intersection numbers
    Intersection numbers $E\cdot F\cdot C_{23}$, with $(E,F)\neq(C_{23},C_{23})$, are zero, with the following exceptions:
@@
    Intersection numbers $E\cdot F\cdot D_{01}$, with $(E,F)\neq(D_{01},D_{01})$, are zero, with the following exceptions:
@@
    Intersection numbers $E\cdot F\cdot A_{0}$, with $(E,F)\neq(A_{0},A_{0})$, are zero, with the following exceptions:
@@ 7.3. The anticanonical divisor of $X$
As a consequence of our study of the geometry of the boundary divisors in $X$ and their intersections, we will show in this section that the anticanonical divisor of $X$ is nef and big; in particular, $X$ is a Mori Dream Space. Let us start with the following straightforward observation:
@@ 8.1. The cone of curves of $X$
We can thus assume that all the fibers of the contraction $\varphi_{R}$ are one dimensional. Then, by the fiber locus inequality (see [32, Theorem 1.1]), $\varphi_{R}$ is a contraction of fiber type. By the assumption on the minimality of $\Gamma$ in $R$, the contraction $\varphi_{R}$ is a ${\mathbb{P}}^{1}$-bundle. We conclude observing that, in this situation, every boundary divisor whose intersection number with $\Gamma$ is zero contains fibers of $\varphi_{R}$. ∎
@@ 8.3. $W$-invariant contractions
$\begin{array}[]{l}M_{1}=\langle L_{1},A_{0}+C_{23}+D_{01},A_{1}+C_{23}+D_{01},A_{2}+C_{01}+D_{23},A_{3}+C_{01}+D_{23}\rangle\\ M^{\prime}_{1}=\langle L_{1},B_{0}+C_{01}+D_{01},B_{1}+C_{01}+D_{01},B_{2}+C_{23}+D_{23},B_{3}+C_{23}+D_{23}\rangle\end{array}$
@@
$L_{2}=A_{2}+A_{3}+2C_{01}+C_{23}+D_{01}+2D_{23},\quad L^{\prime}_{2}=B_{0}+B_{1}+2C_{01}+C_{23}+2D_{01}+D_{23}.$

### v2

@@ 1. Introduction
On the other hand, Kapranov, Białynicki-Birula, Sommese, et al. (see [21], [5]) considered the problem of constructing an intrinsic notion of quotients by group actions, and proposed the closely related concepts of Chow and Hilbert quotients of a variety. The Chow quotient of a variety $F$ by a torus action is the subvariety of the Chow scheme of $F$ parametrizing limits of closures of general orbits.
@@
[…]
Rational homogeneous varieties are defined as quotients of semisimple algebraic groups by parabolic subgroups, and occur in families, classified by the underlying semisimple group $G$. For a fixed $G$, all rational homogeneous $G$-varieties arise as smooth fiber type contractions of the associated complete flag manifold, which is the quotient of $G$ by a Borel subgroup $B$. This is the rational homogeneous variety that is “closest” to $G$, and in fact it has been shown that the homogeneous structure of $G/B$ can be reconstructed out of geometric properties of its elementary contractions (cf. [29], [30, Appendix A]). It makes then sense to start the research by looking at Chow quotients of complete flag varieties, from which the rest of quotients of rational homogeneous varieties can be obtained by means of contractions (see Section 8.3).
We note here that the geometry of these quotients is substantially different from the varieties defining them. First of all, they can be singular, as one can check for the complete flag associated to the group $\mathop{\rm Sp}\nolimits(4)$ (cf. [6]). Then, while the geometry of a complete flag $F=G/B$ is determined by the action of the group $G$, the Chow quotient $X$ of $F$ is naturally endowed with the action of the Weyl group $W$ of $G$. Moreover, while all the contractions of $F$ are smooth fibrations, its quotient $X$ should enjoy a frantic birational life (as the results in [28, 25] already suggest), which makes it an interesting object in the framework of algebraic geometry.
With this in mind, the Chow quotient of the complete flag variety of linear subspaces of ${\mathbb{C}}^{4}$ by its natural torus action, which we explicitly describe in this work, is of particular interest for the following reasons: while being the first nontrivial case of the family of flags of type $\rm A$} – since the Chow quotient of complete flags of ${\mathbb{C}}^{3}$ is ${\mathbb{P}}^{1}$– this Chow quotient is smooth, and its birational geometry, although very rich, can still be explicitly illustrated thanks to a description of the variety as a sequence of blow-ups of ${\mathbb{P}}^{3}$ and to the understanding of its automorphism group. Our main result is the following:
@@
Let $F$ be the complete flag variety of vector subspaces of ${\mathbb{C}}^{4}$, endowed with the natural action of the complex torus $H\subset\mathop{\rm PGL}\nolimits(4)$ of homothety classes of diagonal matrices. Then the Chow quotient $X$ of $F$ by the action of $H$ is a smooth rational weak Fano threefold of Picard number $12$. In particular it is a Mori Dream Space.
We remark that in [13, Theorem 1.3] it was shown, using different methods, that this Chow quotient $X$ is smooth (see Section 9). A large part of this paper is devoted to an explicit, projective-geometric description of $X$ (to out best knowledge, unknown) and of its birational properties. We show that $X$ admits the action of an index two extension of the Weyl group, namely, the octahedral symmetry group $S_{4}\rtimes{\mathbb{Z}}/2{\mathbb{Z}}$. This action is then inherited by the space of numerical class of divisors ${\mathop{\rm N^{1}}\nolimits}(X)$, and by the nef and effective cones of $X$, which we describe.
@@
Sections 7, 8 of the paper focus on the birational geometry of the variety $X$. After studying its intersection theory in terms of boundary divisors – which generate $\mathop{\rm Pic}\nolimits(X)$ – we prove that $X$ is a smooth weak Fano manifold, whose anticanonical model has degree $12$ in ${\mathbb{P}}^{13}$ (cf. Section 7.3); in particular this concludes the proof of Theorem 1.1.
As a consequence $X$ is a Mori Dream Space, therefore its Mori, Nef and Effective cones are polyhedral. We compute these cones in Section 8, and study some of the contractions of $X$. In particular, we identify the Chow quotients of partial flag varieties of ${\mathbb{C}}^{4}$ as targets of contractions of $X$, which we describe explicitly. Moreover, our description of the Mori cone allows us to prove that the group of automorphisms of $X$ is precisely $W(\tau)$.
We conclude the paper with a section containing some final remarks on the results obtained, their relation with other results in the literature, and expectations on further developments in the problem.
@@
Acknowledgements. We would like to thank J. Hong for presenting this problem to us, as well as J. Wiśniewski, M. Donten-Bury and E. Lee for interesting conversations and helpful references, and X. Wu for stimulating discussions on the relation of our work with the theory of matroids.
@@ 2.2. Torus actions and associated quotients
Consider again the above setting. Following [2, §2], we can describe a stronger relation among $\mathcal{C}\mkern-1.0muZ$ and the GIT quotients of $Z$. Let $u,v\in\Gamma$ be such that $Z^{ss}_{v}\subset Z^{ss}_{u}$. Taking quotients we obtain a dominant projective morphism $f_{v,u}:\mathcal{G}\!Z_{v}\to\mathcal{G}\!Z_{u}$. One may show that the collection of GIT quotients $\mathcal{G}\!Z_{v}$, for $v\in\Gamma$, together with the morphisms $f_{v,u}$, for $u\in\Gamma$ such that $Z^{ss}_{v}\subset Z^{ss}_{u}$, form an inverse system. The inverse limit of this system might be reducible (cf. [22, p. 654], [32, Counterexample 1.11]), but there exists a unique irreducible component $\overline{\mathcal{L}Z}$ containing the geometric quotient $\mathcal{G}U$. We call $\overline{\mathcal{L}Z}$ the limit quotient of $Z$, and its normalization $\mathcal{L}Z$ the normalized limit quotient of $Z$ by the action of $H$.
@@ 3.1. Fixed flags and their weights
See [27, Propositions 3.6, 3.7] and the references therein. ∎
@@ 4.5. Description of $X_{e}$ as an inverse limit of GIT quotients
  • The difference between these varieties is that the closed sets $\pi(Z_{(24)})\subset X(\Sigma_{+})$, $\pi(Z_{(03)})\subset X(\Sigma_{-})$ are contracted to the point $\pi(Z_{(0234)})$ in $X(\Sigma_{0})$. The common refinement of $\Sigma_{+}$, $\Sigma_{-}$, whose associated toric variety resolves the induced birational map $X(\Sigma_{+})\dashrightarrow X(\Sigma_{-})$, is the quotient fan $\Sigma$. One may check that the map $X(\Sigma_{+})\dashrightarrow X(\Sigma_{-})$ is locally the standard Atiyah flip (cf. [14, Example 11.1.12], [31, § 1.3]), and that we have a Cartesian square:
@@ 5.2. Generators of the group
Given a general element in $F_{e}$, it can be written as the class modulo $B$ of a matrix
$Y=\begin{pmatrix}1&0&0&0\\ y_{11}&1&0&0\\ y_{12}&y_{22}&1&0\\ y_{13}&y_{23}&y_{33}&1\\ \end{pmatrix}$,
In the open set $U$ defined by $y_{11}y_{22}y_{33}\neq 0$ we have a geometric quotient $U\longrightarrow U/H\subset{\mathbb{P}}^{3}$, that can be written as
(5) $\begin{array}[]{l}YB\mapsto[1:y_{1}:y_{2}:y_{3}]\in{\mathbb{P}}^{3},\mbox{ where}\\[4.0pt] y_{1}:=y_{12}y_{11}^{-1}y_{22}^{-1},\quad y_{2}:=y_{23}y_{22}^{-1}y_{33}^{-1},\quad y_{3}:=y_{13}y_{11}^{-1}y_{22}^{-1}y_{33}^{-1}.\end{array}$
The maps $r_{1},r_{2},r_{3},\tau$, considered as birational automorphisms of $F_{e}$, send the element $YB$ to the classes modulo $B$ of the lower triangular factors of the matrices $r_{1}Y$, $r_{2}Y$, $r_{3}Y$, $\tau Y$, respectively. For instance the image of $YB$ via $r_{1}$ is the class of:
(6) $Y=\begin{pmatrix}1&0&0&0\\ y_{11}^{-1}&1&0&0\\ y_{12}y_{11}^{-1}&y_{12}-y_{11}y_{22}&1&0\\ y_{13}y_{11}^{-1}&y_{13}-y_{11}y_{23}&y_{33}&1\\ \end{pmatrix}$,
Quotienting via the map (6) and homogeneizing, the map $r_{1}$ descends to:
$y_{0}:y_{1}:y_{2}:y_{3}]\mapsto[y_{1}-y_{0}:y_{1}:-y_{2}+y_{3}:y_{3}],$
which is a projectivity of ${\mathbb{P}}^{3}$. Analogously, the maps $r_{2},r_{3},\tau$ descend to:
$\begin{array}[]{l}[y_{0}:y_{1}:y_{2}:y_{3}]\mapsto[y_{1}(y_{2}-y_{0}):y_{0}(y_{2}-y_{0}):y_{1}y_{2}:y_{0}y_{3}],\\[4.0pt] [y_{0}:y_{1}:y_{2}:y_{3}]\mapsto[y_{2}:y_{3}:y_{0}:y_{1}],\\[4.0pt] [y_{0}:y_{1}:y_{2}:y_{3}]\mapsto[y_{0}:y_{0}-y_{2}:y_{0}-y_{1}:y_{0}-y_{1}-y_{2}+y_{3}],\end{array}$
respectively. Note that $r_{1},r_{3},\tau$ are projectivities, while $r_{2}$ is a quadratica Cremona transformation. Finally, we consider the change of homogeneous coordinates, that leaves the above maps in a nicer form:
(7) $\begin{pmatrix}x_{0}\\ x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}1&0&0&0\\ 1&-1&0&0\\ 0&0&1&0\\ 0&0&1&-1\end{pmatrix}\begin{pmatrix}y_{0}\\ y_{1}\\ y_{2}\\ y_{3}\end{pmatrix}$
The outcome of the whole procedure is the following:
@@ 7.2. Intersection numbers
    Intersection numbers $E\cdot F\cdot C_{23}$, with $E,F\neq C_{23}$, are zero, with the following exceptions:
@@
    Intersection numbers $E\cdot F\cdot D_{01}$, with $E,F\neq D_{01}$, are zero, with the exceptions:
@@
    Intersection numbers $E\cdot F\cdot A_{0}$, with $E,F\neq A_{0}$, are zero, with the following exceptions:
@@ 7.3. The anticanonical divisor of $X$
As a consequence of our study of the geometry of the boundary divisors in $X$ and their intersections, we will show in this section that the anticanonical divisor of $X$ is nef and big; in particular, $X$ is a Mori Dream Space. Moreover, $K_{X}$ is globally generated, and its anticanonical model is a $3$-fold of degree $12$ in ${\mathbb{P}}^{13}$. Let us start with the following:
@@ 8.1. The cone of curves of $X$
We can thus assume that all the fibers of the contraction $\varphi_{R}$ are one dimensional. Then, by the fiber locus inequality (see [34, Theorem 1.1]), $\varphi_{R}$ is a contraction of fiber type. By the assumption on the minimality of $\Gamma$ in $R$, the contraction $\varphi_{R}$ is a ${\mathbb{P}}^{1}$-bundle. We conclude observing that, in this situation, every boundary divisor whose intersection number with $\Gamma$ is zero contains fibers of $\varphi_{R}$. ∎
@@
Thanks to the explicit description of the Mori cone of $X$, we can prove the following:
Corollary 8.6.
The automorphism group of $X$ coincides with the octahedral symmetry group $W(\tau)$.
Proof.
By Corollary 6.6 we know that $W(\tau)\subseteq\mathop{\rm Aut}\nolimits(X)$. Notice that an automorphism $f$ of $X$ induces an automorphism $f_{*}$ of ${\mathop{\rm N_{1}}\nolimits}(X)$; we denote by $W^{\prime}$ the image of $\mathop{\rm Aut}\nolimits(X)$ in $\mathop{\rm GL}\nolimits({\mathop{\rm N_{1}}\nolimits}(X))$, and note that, by the projection formula, $\overline{\mathop{\rm NE}\nolimits}(X)$ is $W^{\prime}$-invariant.
We claim first that $\mathop{\rm Aut}\nolimits(X)\to\mathop{\rm GL}\nolimits({\mathop{\rm N_{1}}\nolimits}(X))$ is injective. Since $\overline{\mathop{\rm NE}\nolimits}(X)\subset{\mathop{\rm N_{1}}\nolimits}(X)$ is nondegenerate, it suffices to show that if $f_{*}\in W^{\prime}$ fixes $\overline{\mathop{\rm NE}\nolimits}(X)$ pointwise, then $f$ is the identity map. If $f_{*}$ fixes $\overline{\mathop{\rm NE}\nolimits}(X)$, then it fixes all the extremal faces, and thus all the contractions of $X$. In particular, $f$ fixes the contraction $q:X\to{\mathbb{P}}^{3}$ and descends to an automorphism of ${\mathbb{P}}^{3}$, fixing the images of all the boundary divisors; we conclude that $f$ is the identity. Finally, a computation on SageMath shows that the combinatorial automorphism group of the polyhedral cone of $\overline{\mathop{\rm NE}\nolimits}(X)$, which contains $W^{\prime}$, has order $48$. We conclude that $W(\tau)\simeq\mathop{\rm Aut}\nolimits(X)\simeq W^{\prime}$. ∎
@@ 8.3. $W$-invariant contractions
$\begin{array}[]{l}M_{1}=\langle L_{1},A_{0}+C_{23}+D_{01},A_{1}+C_{23}+D_{01},A_{2}+C_{01}+D_{23},A_{3}+C_{01}+D_{23}\rangle,\\ M^{\prime}_{1}=\langle L_{1},B_{0}+C_{01}+D_{01},B_{1}+C_{01}+D_{01},B_{2}+C_{23}+D_{23},B_{3}+C_{23}+D_{23}\rangle.\end{array}$
@@
$\begin{array}[]{l}L_{2}=A_{2}+A_{3}+2C_{01}+C_{23}+D_{01}+2D_{23},\\ L^{\prime}_{2}=B_{0}+B_{1}+2C_{01}+C_{23}+2D_{01}+D_{23}.\end{array}$
@@ 8.4. The effective cone of $X$
9. Final remarks
We have presented in this paper a projective-geometric description of the normalized Chow quotient $X$ of the complete flag $F$ of type ${\rm A}_{3}$ by the action of a Cartan torus $H$, showing how to obtain it upon the space ${\mathbb{P}}^{3}$ (which is a geometric quotient of $F$) by means of smooth blowups. The main idea we have used is the fact that $X$ is an inverse limit of combinatorial quotients $X_{\sigma}$ of $H$-invariant open subsets of $F_{\sigma}\subset F$, $\sigma\in W$. The way in which this inverse limit is constructed is determined by the action on the system of combinatorial quotients $X_{\sigma}$ of the octahedral group $W(\tau)$, which can be seen as a group of birational automorphisms of ${\mathbb{P}}^{3}$ and, finally, as the group of automorphisms of $X$. Our approach allows us to describe in detail the rich birational geometry of $X$, highlighting the role of the action of $W(\tau)$, which was one of our main goals.
As a by-product, our study provides also a description of $X$ as the union of an open subset (parametrizing the closures of the set of $H$-orbits that contain every $H$-fixed point of the action) and $20$ boundary divisors, parametrizing reducible $H$-invariant cycles (that are degenerations of the closure of the general $H$-orbit). The boundary divisors are the strict transforms of boundary divisors in the combinatorial quotients $X_{\sigma}$, which are toric varieties. Roughly speaking, the boundary divisors in $X_{\sigma}$ tell us how the closure of the general $H$-orbit degenerates as a union of toric varieties locally around the fixed point $\sigma B$.
In this way, this information (together with the intersection theory of boundary divisors, that we use in the study of the birational geometry of $X$) may be used also to describe the moment polytopes of the irreducible components of all the cycles parametrized by $X$. Every reducible cycle is known to correspond to a polytope subdivision of the permutohedron $P$.
We remark that certain properties of the Chow quotient $X$, such as its smoothness, and its description as a log crepant resolution of the log canonical compactification of a certain subset of $F$, were shown in [13]. Their approach, using initial degenerations of the flag variety and matroids, produces in particular a description (see [13, Tables B]), written in terms of flag matroids, of the possible polytope subdivisions of $P$. As a corollary of our study of the boundary divisors, we notice that such list can also be understood as follows:
  • The three coarsest matroidal subdivisions of the permutohedron contained in Table B.2 correspond to the cycles parametrized by general points in the divisors of type $A/B$, $C$ and $D$, respectively.
  • The six intermediate matroidal subdivisions in Table B.3 correspond to the cycles parametrized by general points in (images via elements of $W(\tau)$ of) the curves $A_{0}\mathbin{\mkern-2.5mu*\mkern-2.5mu}C_{23}$, $A_{0}\mathbin{\mkern-2.5mu*\mkern-2.5mu}B_{1}$, $A_{0}\mathbin{\mkern-2.5mu*\mkern-2.5mu}B_{0}$, $A_{0}\mathbin{\mkern-2.5mu*\mkern-2.5mu}D_{01}$, $C_{01}\mathbin{\mkern-2.5mu*\mkern-2.5mu}C_{23}$, and $C_{01}\mathbin{\mkern-2.5mu*\mkern-2.5mu}D_{01}$, respectively.
  • The five finest subdivisions of Table B.4 correspond to points in the Chow quotient that are intersections of three boundary divisors. These points are precisely the $W(\tau)$-orbits of the intersection points represented in Figure 8. We have represented in Figure 9 the subdivision corresponding to points of the form $A_{i}\mathbin{\mkern-2.5mu*\mkern-2.5mu}B_{j}\mathbin{\mkern-2.5mu*\mkern-2.5mu}D_{ij}$.
Figure 9. Subdivisions of $P$ over the points $A_{i}\mathbin{\mkern-2.5mu*\mkern-2.5mu}B_{j}\mathbin{\mkern-2.5mu*\mkern-2.5mu}D_{ij}$.
We conclude by noting that further computer aided experiments show that obtaining detailed descriptions of the Chow quotient of other complete flags $F=G/B$ might be extremely challenging: first of all, a general description of the combinatorial quotients $X_{\sigma}$ is still unknown, even in the case of ${\rm A}_{n}$ (a direct computation of the quotient fan can be fulfilled in a reasonable time in the case of ${\rm A}_{4}$, but the case of ${\rm A}_{5}$ is already too computationally demanding); then the simplest geometric quotient of $F$ may not be the projective space (as in the case ${\rm A}_{3}$), but a weighted projective space, which makes the study of the action of the Weyl group on quotients of $F$ more complicated; last but not least, the description of the Chow quotient of the complete flag of type ${\rm C}_{2}$ (cf. [6]) suggests that Chow quotients of flags may be rather singular, in general. However, we still expect that our presentation of the Chow quotient as a limit of combinatorial quotients (cf. § 4.5) may be used to describe some of its general properties.